# UniMS-RAG: A Unified Multi-source Retrieval-Augmented Generation for Personalized Dialogue Systems

### v1

@@ 4. Method
Figure 2. Our proposed method UniMS-RAG, where three optimization tasks are carefully designed: 1) Knowledge Source Selection; 2) Relevance Score Prediction; and 3) Response Generation.
4.1. UniMS-RAG Framework
Building upon the foundation of SAFARI (Wang et al., 2023a), we propose an innovative methodology termed UniMS-RAG, where $U$ signifies the unification of the training process for planner, retriever and reader, as well as the integration of diverse tasks into a singular comprehensive framework. Recognizing the potential of large language models (LLMs) in orchestrating the utilization of external sources of knowledge, as indicated by recent works, UniMS-RAG extends the capabilities of LLMs to seamlessly connect disparate sources within the context of personalized knowledge-grounded dialogues. This integration streamlines the traditionally separated tasks of retriever and reader training, enabling to adaptively retrieve evidences and evaluate the relevance scores in a unified manner.
@@ 4.1. UniMS-RAG Framework
4.1.1. Planning
First of all, to incorporate the cases which does not require any sources of external knowledge, we define several additional indicate tokens, corresponding to different sources, including the NULL token which signifies that there is no requirement for any external knowledge as shown in Table 1:
(1) $\mathcal{M}:\boldsymbol{c}\rightarrow\texttt{NULL},$
Where $\mathcal{M}$ is parameterized by LLMs. Secondly, there are two different situations between these multiple knowledge sources: 1) the $K_{1},K_{2},...,K_{i}$ in $\boldsymbol{K}$ is independent, which means there is no interdependent relationship between them; and 2) Some or even all of they are not independent, for example, the results obtained from $K_{1}$ may be contingent on the outcomes derived from $K_{2}$ and potentially other sources. These two situations cater to different applications. In the first scenario, the independence between knowledge sources offers practical utility for tasks where autonomy and isolation of information are paramount, such as user persona and system persona, which are two independent knowledge sources. On the other hand, in the second scenario, where interdependencies appear, the model accommodates tasks that demand a nuanced consideration of the relationships between different knowledge sources. For example, dependencies between user memory or persona and document introduce a layer of complexity, reflecting real-world scenarios where the fact that age, hobby, education, and life experience of the user have a major effect on his or her personal preference over external document knowledge (Fu et al., 2022).
Table 1. Three types of special tokens used in UniMS-RAG, in which the former two stands for acting tokens and evaluation tokens, respectively. Each type uses several tokens to represent its output values.
@@ 4.1.1. Planning
(2) $\mathcal{M}:\boldsymbol{c}\rightarrow K_{i},K_{j},...,K_{n},$
Then we strictly follow the outputed order to retrieve evidences from corresponding source of knowledge. To offer the flexibility and scalability to plug in an arbitrary number of sources, we can add $K_{i},...,K_{n}$ and NULL as special tokens into the vocabulary of LLMs as special tokens, and expand the set of tokens on the fly following Hao et al. (Hao et al., 2023). Besides that, we add other special tokens to indicate the different parts of the input, i.e., [SOURCE] and [EOS] to indicate the start and end positions of sources. In this way, LLM can model the dependency between different sources and learn when and how to call certain sources.
4.1.2. Retrieval
According to the output of the last planning step, there are two cases in this step: (1) the response does not need any external sources of knowledge, and the dialogue system can skip this step; and (2) the response needs multiple sources of knowledge, and it strictly follows the output source order to retrieve top-n related evidence $k_{i}^{*}$ for the $i_{th}$ source of knowledge according to the dialogue context $\boldsymbol{c}$. If there is a dependency here, it will use preceding retrieved results $k_{j}^{*}$ in the planned execution order as a filter. Specifically, assuming the output order is PERSONA, DOCUMENTS in the planning step for a persona-consistent dialogue system, we first retrieve top-1 result $p^{*}$ from PERSONA, and then we retrieve $k^{*}$ from DOCUMENTS according to $\boldsymbol{c}$ and $p^{*}$ (Figure 0(b)). If there is no dependency, it simply retrieves corresponding evidences source by source (Figure 0(a)).
To joint train the retriever and reader in a unified manner, we first reformulate the traditional classification task (relevant or irrelevant) for retrieval into a generation task. Then we define several similarity tokens as shown in Table 1, ranging from 0 to 1 with the internal as 0.1. In this way, we enforce the UniMS-RAG to predict the similarity score by generating corresponding tokens. To get the supervised label of similarity scores, there are two different ways: 1) the similarity scores are generated by independent trained retriever such as DPR (Karpukhin et al., 2020); 2) the similarity scores are generated by prompting LLMs (Zhu et al., 2023; Sun et al., 2023b). The details can be found in § 4.2. Furthermore, we design a evidence attention mask mechanism to mask unrelated evidence when predicting the similarity score for current evidence, aiming to reduce unnecessary noises in the context as shown in Figure 2.
(3) $\mathcal{M}:\boldsymbol{c},K_{i},e_{j}\rightarrow sim\in\{0.1,0.2,...,1.0\},$
@@ 4.1.3. Generation
4.2. Training Stage
In this section, we first introduce how to acquire the soft signals of these relevance scores and then present the global training objectives of our proposed UniMS-RAG.
4.2.1. Relevance Scores Acquisition
There are two ways to get the soft labels to indicate the relevance between the dialogue context and evidence: 1) using a fine-tuned DPR (Karpukhin et al., 2020) model; and 2) directly prompting the LLMs (i.e, ChatGPT) (Sun et al., 2023a; Ma et al., 2023). Here we present the details of these methods one by one.
(1) DPR as a Retriever: Following previous works (Karpukhin et al., 2020), we can train an off-the-shelf retriever independently. Specifically, we first build our own finetuning dataset by regarding (context, related_evidence) as the positive and (context, unrelated_evidence) as the negatives. Then we apply the following negative log likelihood (NLL) loss to learn the retrieval-oriented representations following (Yu et al., 2021):
(5) $\mathcal{L}_{retriever}=-log\frac{e^{sim(C_{i},d^{+})}}{e^{sim(C_{i},d^{+})}+\sum_{j\in\mathcal{B^{-}}}{e^{sim(C_{i},d^{-})}}},$
(6) $sim(C,d)=cos(C,d))$
@@ 4.2.1. Relevance Scores Acquisition
(2) LLMs as a Retriever: Besides these supervised training methods for retriever, we can utilize some off-the-shelf methods to get the similarity score between the dialogue context and the evidence, including some sparse retriever such as TF-IDF (Sparck Jones, 1972) and BM25 (Robertson et al., 2009), or simply hard label by assigning a similarity score to the evidence based on whether it is used (set score as 1) or unused (set score as 0). Inspired by recent studies which simply using LLMs as search engine to predict the similarity relationship between query and evidence (Sun et al., 2023b; Ma et al., 2023), we also choose to prompt the LLMs (i.e, ChatGPT) to predict the similarity score by feeding a dialogue context and several evidences in a zero-shot manner. The prompts are shown in Figure 3.
Figure 3. The instructions for zero-shot retriever used to predict similarity score using off-the-shelf LLMs. The grey and yellow blocks indicate the inputs and outputs of the model.
During the inference stage, three types of retrievers can be employed: 1) DPR, 2) LLMs, and 3) UniMS-RAG. It is worth noting that UniMS-RAG can be considered a distillation derived from either DPR or LLMs.
4.2.2. Training Objectives
The whole training objective of UniMS-RAG can be defined as follows:
(7) $\mathcal{L}=\mathcal{L}_{source}+\mathcal{L}_{sim}+\mathcal{L}_{response}$
Where the $L_{source},L_{sim},L_{response}$ indicate planning loss, relevance predication loss and the final response loss. By assigning different tokens different roles, our UniMS-RAG can serve as the planner (determine whether or not use external sources of knowledge and then use which source), retriever, and reader in a unified manner.
4.3. Inference Stage
@@
6:  S = $S_{nli}$ $S_{ce}$ // element-wise product inspired by $p(r|c)=p(e,c)p(r|c,e)$
@@
5. Experimental Setups
5.1. Research Questions
@@
5.2. Datasets
Table 2. Statistics of DuLeMon and KBP Datasets
Dataset DuLeMon KBP
#Dialogues 3011 2477
Train/Val/Test 19437/2407/2416 9821/1227/1229
#Utterances 24554 48522
@@ 5.2. Datasets
Source Types NULL, User-Per, Bot-Per NULL, Persona, Document
Resp w/ source $\approx$ 49% $\approx$ 85%
@@
We consider two different situations between different knowledge sources, corresponding to two publicly available personalized dialogue datasets: DuLeMon [^3] https://github.com/PaddlePaddle/Research/tree/master/NLP/ACL2022-DuLeMon (Xu et al., 2022a) and KBP  [^4] https://github.com/ruleGreen/SAFARI (Wang et al., 2023a). The dataset statistics are presented in Table 2. We adopt the same train/val/test split in the original datasets.
@@
  • DuLeMon (Xu et al., 2022a) is the latest open-domain dialogue dataset with long-term persona memory in which a response is grounded on persona information that occurred in historical sessions, leading to better dialogue engagement. There are two versions of DuLemon: Self and Both, where the persona information comes from only self side (user side) or both side (both user and chatbot side). We choose Both versions to consider the independent relationship between these two sources of persona information: User side (User-Per) and Chatbot Side (Bot-Per).
@@
  • Knowledge Behind Persona (a.k.a., KBP) (Wang et al., 2023a) is a dialogue dataset, in which the response is grounded on both the persona (Persona) and its corresponding implicit knowledge (Document). We choose it to consider the interdependent relationship between different sources of knowledge. It is worth noting that both of these two datasets contain cases which do not require any external source of knowledge, denoting as NULL.
5.3. Baselines and Evaluation Metrics
5.3.1. Knowledge Source Selection
We first include two methods to select different sources of knowledge for comparisons:
@@ 5.3.1. Knowledge Source Selection
  • BERT (Devlin et al., 2019) We utilize BERT as the backbone to train a classifier, which determines the required sources. We formulate this task as a multi-label classification task to determine whether each source is required or not required. If none of them is required, then the used source is NULL. We choose different thresholds for different datasets according to the performance at validation dataset  [^5] In detail, we set the threshold as 0.3 in kBP and 0.5 in DuLeMon datasets, respectively..
@@
  • SAFARI (Wang et al., 2023a) is the first work to formulate the source planning task as a sequence generation task, regarding different sources of knowledge as different tokens in vocabulary at LLMs.
Following previous works (Wang et al., 2023a), we adopt F1 as automatic evaluation metrics.
5.3.2. Knowledge Retrieval
We compare the performance of retrieval with different types of retrievers:
  • BM25 (Robertson et al., 2009) is a type of sparse retriever, which takes into account term frequencies and inverse document frequency (TF-IDF) to score the relevance of documents to a query.
  • RocketQAv2 (Ren et al., 2021) is a type of dense retriever, which is a unified list-wise training approach for both the retriever and the re-ranker.
  • DPR (Karpukhin et al., 2020) is a method that leverages dense vector representations for passages in a collection. It uses pre-trained language models to encode passages and queries into dense vectors independently, allowing for a more nuanced understanding of the content.
  • ChatGPT [^6] https://openai.com/chatgpt We utilize the same instruction as shown in Figure 3 to prompt gpt-3.5-turbo-1106 to retrieve top-n evidence from corresponding source of knowledge.
@@ 5.3.2. Knowledge Retrieval
We have chosen Recall@1 as our primary evaluation metric. This selection is motivated by the predominant use case scenario observed in the original datasets, where a single evidence from each knowledge source is typically sufficient.
5.3.3. Response Generation
We mainly compare two methods due to limited work focus on our target problem:
  • FoCus (Jang et al., 2022b) aims to minimize the negative log-likelihood of language modeling and sub-tasks: persona grounding and knowledge grounding. It uses either the sigmoid or softmax functions to select evidence by concatenating evidence and dialogue context, and then generate the final response. We do not report the performance on DuLeMon dataset since most of responses in this dataset do not require external sources.
  • SAFARI (Wang et al., 2023a) incorporates all retrieved evidences with corresponding source signals to generate the final responses, without considering the relevance between different evidences and dialogue context explicitly. We report the performance of supervised and unsupervised SAFARI, following Wang et al. (2023a).
@@ 5.3.3. Response Generation
Following Wang et al. (2023a) and Xu et al. (2022a), we choose BLEU, Rouge-L to evaluate the gap between generated response with the ground truth response in the original datasets. Besides that, we select Persona Consistency (P.C) and Knowledge Consistency (K.C) to evaluate the consistency score using our finetuned NLI models (Madotto et al., 2019). Specifically, there are four cases during experiments as shown in Eq 11: 1). there exists ground-truth grounding $g$ for the response $r$ in the test set and the planning also decides to retrieve ground knowledge $g^{\prime}$ from corresponding sources; 2). there is $g$ for $r$, but the planning decides not to use $g^{\prime}$; 3). there is no grounding $g$ for $r$, while the planning decides to use $g^{\prime}$; 4). there exists no $g$, and the planning step decides not to use $g^{\prime}$ either. We only calculate the NLI score of $(g,r)$ using finetuned NLI model in the first case of Eq 11. With this definition, we can get the calibrated score P.C [^7] It is important to note we merge the two sources of persona information (User-Per, Bot-Per) in DuLeMon into one unified P.C score. and K.C:
(8) $\textbf{NLI}(r,g)=\begin{cases}1,&\text{if $r$ entails $g$}\\ 0,&\text{if $r$ does not entail $g$}\\ \end{cases}$

### v2

@@ 4 Method
Figure 2: Our proposed method UniMS-RAG, where three optimization tasks are carefully designed: 1) Knowledge Source Selection; 2) Relevance Score Prediction; and 3) Response Generation. We use orange to indicate acting tokens and blue to indicate evaluation tokens. It is worth noting we have all labels during training to optimize these three sub-tasks in a teacher-forcing way.
4.1 UniMS-RAG
Inspired by previous work [11], we propose an innovative methodology termed UniMS-RAG, where $U$ signifies the unification of the training process for planner, retriever and reader, as well as the integration of diverse tasks into a singular comprehensive framework. Recognizing the potential of large language models (LLMs) in orchestrating the utilization of external sources of knowledge, as indicated by recent works, UniMS-RAG extends the capabilities of LLMs to seamlessly connect disparate sources within the context of personalized knowledge-grounded dialogues. This integration streamlines the traditionally separated tasks of retriever and reader training, enabling to adaptively retrieve evidences and evaluate the relevance scores in a unified manner.
@@ 4.1 UniMS-RAG
4.1.1 Planning
First of all, to incorporate the cases which does not require any sources of external knowledge, we define several additional indicate tokens, corresponding to different sources, including the NULL token which signifies that there is no requirement for any external knowledge as shown in Table I:
$\mathcal{M}:\boldsymbol{c}\rightarrow\texttt{NULL},$ (1)
where $\mathcal{M}$ is parameterized by LLMs. Secondly, there are two different situations between these multiple knowledge sources: 1) the $K_{1},K_{2},...,K_{i}$ in $\boldsymbol{K}$ is independent, which means there is no interdependent relationship between them; and 2) Some or even all of they are not independent, for example, the results obtained from $K_{1}$ may be contingent on the outcomes derived from $K_{2}$ and potentially other sources. These two situations cater to different applications. In the first scenario, the independence between knowledge sources offers practical utility for tasks where autonomy and isolation of information are paramount, such as user persona and system persona, which are two independent knowledge sources. On the other hand, in the second scenario, where interdependencies appear, the model accommodates tasks that demand a nuanced consideration of the relationships between different knowledge sources. For example, dependencies between user memory or persona and document introduce a layer of complexity, reflecting real-world scenarios where the fact that age, hobby, education, and life experience of the user have a major effect on his or her personal preference over external document knowledge [28].
TABLE I: Three types of special tokens used in UniMS-RAG, in which the former two stands for acting tokens and evaluation tokens, respectively. Each type uses several tokens to represent its output values.
@@ 4.1.1 Planning
$\mathcal{M}:\boldsymbol{c}\rightarrow K_{i},K_{j},...,K_{n},$ (2)
Then we strictly follow the outputed order to retrieve evidences from corresponding source of knowledge. To offer the flexibility and scalability to plug in an arbitrary number of sources, we can add $K_{i},...,K_{n}$ and NULL as special tokens into the vocabulary of LLMs as special tokens, and expand the set of tokens on the fly following Hao et al. [49]. Besides that, we add other special tokens to indicate the different parts of the input, i.e., [SOURCE] and [EOS] to indicate the start and end positions of sources. In this way, LLM can model the dependency between different sources and learn when and how to call certain sources.
4.1.2 Retrieval
According to the output of the last planning step, there are two cases in this step: (1) the response does not need any external sources of knowledge, and the dialogue system can skip this step; and (2) the response needs multiple sources of knowledge, and it strictly follows the output source order to retrieve top-n related evidence $k_{i}^{*}$ for the $i_{th}$ source of knowledge according to the dialogue context $\boldsymbol{c}$. If there is a dependency here, it will use preceding retrieved results $k_{j}^{*}$ in the planned execution order as a filter. Specifically, assuming the output order is PERSONA, DOCUMENTS in the planning step for a persona-consistent dialogue system, we first retrieve top-1 result $p^{*}$ from PERSONA, and then we retrieve $k^{*}$ from DOCUMENTS according to $\boldsymbol{c}$ and $p^{*}$. If there is no dependency, it simply retrieves corresponding evidences source by source (as shown in Figure 1).
To joint train the retriever and reader in a unified manner, we first reformulate the traditional classification task (relevant or irrelevant) for retrieval into a generation task. Then we define several similarity tokens as shown in Table I, ranging from 0 to 1 with the internal as 0.1. In this way, we enforce the UniMS-RAG to predict the similarity score by generating corresponding tokens. To get the supervised label of similarity scores, there are two different ways: 1) the similarity scores are generated by independent trained retriever such as DPR [42]; 2) the similarity scores are generated by prompting LLMs [45]. The details can be found in § 4.2. Furthermore, we design a evidence attention mask mechanism to mask unrelated evidence when predicting the similarity score for current evidence, aiming to reduce unnecessary noises in the context as shown in Figure 2.
$\mathcal{M}:\boldsymbol{c},K_{i},e_{j}\rightarrow sim\in\{0.1,0.2,...,1.0\},$ (3)
@@ 4.1.3 Generation
[…]
. We use [EVIDENCE] and [EOE] to represent the start and end positions of the retrieved evidences. The $[Sim_{i}]$ stands for the similarity score of $i_{th}$ evidence, calculated using the retriever (§4.2).
4.2 General Framework
Besides the core model, we also value the general framework such as different ways to collect data, unique training and inference strategies. Thus, we first introduce different ways to collect relevance score labels, and then present the training and inference design (Figure 3).
4.2.1 Relevance Scores Acquisition
There are two different methods to get the similarity labels: 1) prompt-based method. which directly prompt the LLMs to assign the similarity scores given dialogue context and evidences [46, 48]; 2) classification-based method. which requires a fine-tuned retriever such as DPR to calculate the similarity score given the same input. In latter method, the evidence is feed one by one with the dialogue context unchanged.
Prompt-based method. We can utilize some off-the-shelf methods to get the similarity score between the dialogue context and the evidence, including some sparse retriever such as TF-IDF [50] and BM25 [40], or simply hard label by assigning a similarity score to the evidence based on whether it is used (set score as 1) or unused (set score as 0). Inspired by recent studies which simply using LLMs as search engine to predict the similarity relationship between query and evidence [48], we choose to prompt the LLMs (i.e, ChatGPT) to predict the similarity score by feeding a dialogue context and several evidences in a zero-shot manner. The prompts are shown in Figure 4.
Figure 3: The general framework to utilize UniMS-RAG, including 1) relevance score acquisition; 2) training stage; and 3) inference stage.
Figure 4: The instructions for zero-shot retriever used to predict similarity score using off-the-shelf LLMs. The grey and yellow blocks indicate the inputs and outputs of the model.
Classification-based method. Following previous works [42], we can train an off-the-shelf retriever using the ground-truth labels. Specifically, we first build our own finetuning dataset by regarding (context, related_evidence) as the positive and (context, unrelated_evidence) as the negatives. The positive and negative samples are annotated in the original dialogue dataset, i.e., which persona or documents is used to generate the response. Then we apply the following negative log likelihood (NLL) loss to learn the retrieval-oriented representations following [51]:
$\mathcal{L}_{retriever}=-log\frac{e^{sim(C_{i},d^{+})}}{e^{sim(C_{i},d^{+})}+\sum_{j\in\mathcal{B^{-}}}{e^{sim(C_{i},d^{-})}}},$ (5)
$sim(C,d)=cos(C,d))$ (6)
@@ 4.2.1 Relevance Scores Acquisition
To avoid computing on the fly during training, we pre-compute relevance scores for all training and validation samples before beginning the model training process. During the inference stage, three types of retrievers can be employed: 1) DPR, 2) LLMs, and 3) UniMS-RAG. It is worth noting that UniMS-RAG can be considered a distillation derived from either DPR or LLMs.
4.2.2 Training Stage
We shuffle the order of evidence during the training. The formation of the input in this way has three advantages. Firstly, the name of the sources indicates the type of results retrieved, which provides more signals to the LLMs. Secondly, it enforces the LLMs to capture the important evidence since it can appear in any position and denoise unrelated evidence. Thirdly, it allows us to train the large language model in a multi-task manner using teacher forcing as introduced in next section. The whole training objective of UniMS-RAG can be defined as follows:
$\mathcal{L}=\mathcal{L}_{source}+\mathcal{L}_{sim}+\mathcal{L}_{response}$ (7)
Where the $L_{source},L_{sim},L_{response}$ indicate planning loss, relevance predication loss and the final response loss. In other words, we only calculate the loss at acting, evaluation and response tokens. By assigning different tokens different roles, our UniMS-RAG can serve as the planner (determine whether or not use external sources of knowledge and then use which source), retriever, and reader in a unified manner.
4.2.3 Inference Stage
@@
6:  S = $S_{nli}$ $S_{ce}$ // element-wise product inspired by $p(r|c)=p(e|c)p(r|c,e)$
@@
5 Experimental Setups
5.1 Research Questions
@@
5.2 Datasets
TABLE II: Statistics of DuLeMon and KBP Datasets
@@ 5.2 Datasets
Dataset DuLeMon KBP #Dialogues 3011 2477 Train/Val/Test 19437/2407/2416 9821/1227/1229 #Utterances 24554 48522 Source Types NULL, User-Per, Bot-Per NULL, Persona, Document Resp w/ source $\approx$ 49% $\approx$ 85%
@@
We consider two different situations between different knowledge sources, corresponding to two publicly available personalized dialogue datasets: DuLeMon [7] and KBP [11]. The dataset statistics are presented in Table II. We adopt the same train/val/test split in the original datasets.
@@
  • DuLeMon [7] is the latest open-domain dialogue dataset with long-term persona memory in which a response is grounded on persona information that occurred in historical sessions, leading to better dialogue engagement. There are two versions of DuLemon: Self and Both, where the persona information comes from only self side (user side) or both side (both user and chatbot side). We choose Both versions to consider the independent relationship between these two sources of persona information: User side (User-Per) and Chatbot Side (Bot-Per).
@@
  • Knowledge Behind Persona (a.k.a., KBP) [11] is a dialogue dataset, in which the response is grounded on both the persona (Persona) and its corresponding implicit knowledge (Document). We choose it to consider the interdependent relationship between different sources of knowledge. It is worth noting that both of these two datasets contain cases which do not require any external source of knowledge, denoting as NULL.
5.3 Baselines and Evaluation Metrics
5.3.1 Knowledge Source Selection
We begin by discussing prompting-based and classification-based methods. This includes presenting several baseline models to select different sources of knowledge for comparison, and we adopt F1 as automatic evaluation metrics, following previous works [11].
  • Prompting-based methods. We mainly utilize zero-shot and in-context learning prompting (ICL) to directly instruct ChatGPT (gpt3.5-turbo-1106) and GPT4o (gpt-4o) to output the required knowledge sources (a.k.a, acting tokens). We use same demonstration for fair comparison.
@@ 5.3.1 Knowledge Source Selection
  • BERT [52] We utilize BERT as the backbone to train a classifier, which determines the required sources. We formulate this task as a multi-label classification task to determine whether each source is required or not required. If none of them is required, then the used source is NULL. We choose different thresholds for different datasets according to the performance at validation dataset  [^2] In detail, we set the threshold as 0.3 in kBP and 0.5 in DuLeMon datasets, respectively..
@@
  • SAFARI [11] is the first work to formulate the source planning task as a sequence generation task, regarding different sources of knowledge as different tokens in vocabulary at LLMs. However, it does not consider relevance scores or iterative refinement. We copy the its best performance for better comparison.
5.3.2 Knowledge Retrieval
@@
We compare the performance of retrieval with different types of retrievers, and we choose Recall@1 as our primary evaluation metric. This selection is motivated by the predominant use case scenario observed in the original datasets, where a single evidence from each knowledge source is typically sufficient.
  • BM25 [40] is a type of sparse retriever, which takes into account term frequencies and inverse document frequency (TF-IDF) to score the relevance of documents to a query.
  • RocketQAv2 [53] is a type of dense retriever, which is a unified list-wise training approach for both the retriever and the re-ranker.
  • DPR [42] is a method that leverages dense vector representations for passages in a collection. It uses pre-trained language models to encode passages and queries into dense vectors independently, allowing for a more nuanced understanding of the content.
  • LLMs. We utilize the same instruction as shown in Figure 4 to prompt gpt-3.5-turbo-1106 and gpt-4o to retrieve top-n evidence from corresponding source of knowledge. Furthermore, we also employ in-context learning (ICL) prompting for more comprehensive comparison  [^3] We randomly sampled three cases to cover all knowledge sources from the original dataset. We do not observe significant improvements with other demonstration selection strategies..
5.3.3 Response Generation
@@
We mainly compare two methods due to limited work focus on our target problem. Regarding the metrics, we choose BLEU, Rouge-L to evaluate the gap between generated response with the ground truth response in the original datasets. Besides that, we select Persona Consistency (P.C) and Knowledge Consistency (K.C) to evaluate the consistency score using our finetuned NLI models [8] with same definition in [11]. It is important to note we merge the two sources of persona information (User-Per, Bot-Per) in DuLeMon into one unified P.C score.
  • FoCus [54] aims to minimize the negative log-likelihood of language modeling and sub-tasks: persona grounding and knowledge grounding. It uses either the sigmoid or softmax functions to select evidence by concatenating evidence and dialogue context, and then generate the final response. We do not report the performance on DuLeMon dataset since most of responses in this dataset do not require external sources.
  • SAFARI [11] incorporates all retrieved evidences with corresponding source signals to generate the final responses, without considering the relevance between different evidences and dialogue context explicitly. We report performance of both supervised and unsupervised SAFARI.
5.4 Implementation Details
UniMS-RAG: We mainly choose ChatGLM-6B [55, 56] as the backbone models during training, we set the batch size as 8, train models with 3 epochs and save the checkpoint with the lowest validation loss. For other hyper-parameter settings, we mainly follow the corresponding official code  [^4] https://github.com/THUDM/ChatGLM-6B. Due to the computation limit, we conduct training with LoRA [57] at one single 3090 GPU, and it costs about 4-6 hours. For the prompting using LLMs, we choose gpt-3.5-turbo-1106 and gpt-4o, and set both the temperature and top p as 0.1 to reduce the randomness of LLMs. We choose only the top-ranked result for the main experiment to ensure a consistent setting, and we additionally investigate the effects of retrieving different numbers of results to evaluate whether UniMS-RAG can effectively filter out unrelated evidence. According to different sources of similarity scores, we have three variants: UniMS-RAG (ChatGPT), UniMS-RAG (GPT-4o) and UniMS-RAG (DPR) where the similarity scores come from ChatGPT, GPT4o and DPR, respectively. At the inference stage, we use self-predicted similarity scores to retrieve the evidence, serving as an indicator. More variants or analyses can be found in §7.2.
Others: We finetune RocketQAv2 and DPR using samples from corresponding dataset by regarding (context, used_persona / document) as the positive and (context, unrelated_persona / document) as the negative. We set epochs as 5 and max sequence length as 512, and mainly follow these codebases [^5] https://github.com/PaddlePaddle/RocketQA, https://github.com/Alibaba-NLP/Multi-CPR/tree/main/retrieval for other parameters. For RocketQAv2, we load the weights of pre-trained model zh_dureader_de_v2 as introduced in the official homepage, which is trained on the largest Chinese QA dataset, and we use 12-layer bert-base-chinese with 110M parameters as backbone model for DPR. We then finetune an NLI model [8] by regarding (ground_persona / document, response) as the positive and randomly sampled (unrelated_persona / document, response) as the negative. We also use bert-base-chinese as the backbone model. We concatenate and encode the ground persona/document $k$ and response $r$ in the form of $\mathrm{[CLS]}k\mathrm{[SEP]}r\mathrm{[SEP]}$, and we train the model to predict whether responses are consistent with corresponding personas or documents. The batch size for fine-tuning is 8. The maximal training epoch is 5, and the max sequence length of the encoder is 512. In the experiments, we use the AdamW optimizer with a learning rate of 2e-5 and an epsilon of 1e-6. We evaluate the NLI model on the KBP test set every 500 iterations during training, and we save the checkpoint with the highest performance on the test set. The fine-tuned NLI model achieves $>95\%$ accuracy for both datasets.
6 Experimental Results
TABLE III: The F1 of different decisions in Planning of different methods under supervised settings. 1) DuLeMon: There are 1686 NULL, 505 USER, 219 BOT and 6 BOTH in the ground planning; 2) KBP: There are 181 NULL, 125 PERSONA and 923 BOTH in the ground planning.